OpenStudy (cenaida):

The expression x^2 - 4 A. Can be factored B. Cannot be factored

OpenStudy (shamil98):

A. Like i said (x+2)(x-2)

OpenStudy (shamil98):

bx get's canceled out because , -2x + 2x = 0

OpenStudy (anonymous):

(x−2)(x+2) = x^2 −4